2. Find the derivative of f(x) = 5(723 – 2x + 1)-3. 3. Differentiate f(x) = (3.2 + 10 – sin x). O e 99+ ASUS VivoBook
Answer #1

Z To find the derivative of f(x) = 5(7x² – 2x+1) 3 f(x) = dx [ 5(x-2x+1)] [(7x?_2x+1)”] dx -3-1 = 5*(-3) (7x3–2x+1) (7x3–2x+